OpenStudy (amanda0618):

What are the vertex and the axis of symmetry of the equation y=2x^2+4x-10 please help and show work!!

OpenStudy (mathmale):

y=2x^2+4x-10 is a quadratic function and has a graph which is a parabola. The coefficients of the x-terms are, in order, 2, 4 and -10. Rewrite this function in "vertex form," \[y=a(x-h)^2+k\], to obtain the coordinates of the vertex: they are (h, k). From the quadratic formula, \[x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a },\]

OpenStudy (mathmale):

we get \[x=-b/(2a) \]

OpenStudy (mathmale):

as the equation of the axis of symmetry. Please show your work, so that someone can give you feedback on what you've already done.

OpenStudy (amanda0618):

what? I'm so confused

jimthompson5910 (jim_thompson5910):

@Amanda0618 do you see how `y=2x^2+4x-10 ` is in the form \(\Large y = ax^2+bx+c\)?

jimthompson5910 (jim_thompson5910):

you may have to write it as y = 2x^2+4x+(-10)

OpenStudy (mathmale):

Compare YOUR quadratic equation with the standard form given you by Jim. Determine through this comparison the values of the coefficients a, b and c. Try finding the equation of the axis of symmetry first. As before, it's x = -b/(2a).

OpenStudy (mathmale):

You have begun with the quadratic equation y = 2x^2+4x+(-10). To find the coordinates of the vertex of the parabolic graph of this function, start by factoring out the coefficient 2: y = 2(x^2 + 2x - 5). You'll need to use the "completing the square" approach to rewrite x^2 + 2x - 5 in the form y = x^2 + 2x + 1 - 1 - 5. Have you used this approach before? If not, I'd suggest you look up "completing the square" and apply what you learn to rewriting x^2 + 2x - 5.

OpenStudy (mathmale):

Here's an example of "completing the square:" Given y = x^2 - 6x + 2, rewrite this in the form y = x^2 - 6x + 9 - 9 + 2, or y = x^2 - 6x + 9 + (-9 + 2). Why? Because x^2 - 6x + 9 is a perfect square. This simplifies to y = (x-3)^2 - 7. The vertex is (3, -7).

OpenStudy (mathmale):

I am trying to get you involved. Please do and then share any work you can do on your own; then I could give you some meaningful feedback. Once again: quadratic equations come in several different forms, all of which you will likely need to learn. y = a(x-h)^2 + k (vertex form) y = 4px^2 (where p is the distance between focus and vertex) y = ax^2 + bx + c ("standard form")

OpenStudy (mathmale):

Welcome to Open Study. I know you're new and are likely exploring what others are typing in. If you want help with a problem of your own, it's best to stick with it for a while, in case someone tries to help you. After five or ten minutes of waiting for a reply from you, I'm ready to move on to helping others.